# An active learning enabled 4th-generation high dimensional committee neural network potential for the simulation of quantum vibrational spectra: Applications to the bulk and the air-water interface with charge transfer and nuclear quantum effects included

### v1

@@ Abstract
Predictive simulation of vibrational spectra of complex condensed phases and interfaces with thousands of degrees of freedom has long been a challenging task of modern condensed matter theory. In this work, fourth-generation high-dimensional committee neural network potentials (4G-HDCNNPs) are developed for the first time, using active learning and query-by-committee, and introduced to the essential nuclear quantum effects (NQEs) as well as conformational entropy and anharmonicities from path integral (PI) molecular dynamics simulations. Using representative bulk water and air-water interface systems, we demonstrate the accuracy of the developed framework in infrared and vibrational sum frequency generation spectral simulations. Specifically, by seamlessly integrating non-local charge transfer interactions from 4G-HDCNNPs with the NQEs from PI methods, our introduced methodology yields accurate infrared spectra using predicted charges from the 4G-HDCNNP architecture without explicit training of dipole moments. Our vibrational sum frequency generation spectra also illustrate high accuracy for the considered air-water interface system. The introduced framework in this work is general and offers a benchmark tool and a simple, practical paradigm for predictive spectral simulations of complex condensed phases and interfaces, free from empirical parameterizations and/or ad hoc fittings.
@@ 1 Introduction
Vibrational spectroscopy, including linear infrared (IR) and non-linear vibrational sum frequency Generation (vSFG) techniques, is an excellent tool for probing complex processes across condensed phases and interfaces.1, 2, 3, 4 To obtain accurate predictive vibrational spectra, free from ad hoc fittings or empirical parameterizations, it has been well-established that robust dynamical methodologies that incorporate anharmonicities and conformational entropies, as well as the essential nuclear quantum effects (NQEs) are needed to be developed and integrated with accurate potential energy surfaces (PESs) that include non-local charge transfer interactions.5, 6 Path integral molecular dynamics (PIMD) offers an accurate and efficient route to include NQEs by representing a quantum particle as a series of identical classical particles, referred to as beads connected via harmonic springs to form a closed loop, referred to as a ring polymer.7, 8, 9 Accurate equilibrium properties that incorporate NQEs can be calculated using PI Monte Carlo (PIMC) or PI molecular dynamics (PIMD) simulations.8 Similarly, approximate real-time PI methods, including thermostatted ring polymer MD (TRPMD),10 partially adiabatic centroid MD (PA-CMD)11, and a myriad of methods branched from them,12, 13, 14, 15, can then be used for the simulation of dynamical properties including vibrational spectra through the calculations of appropriate time correlation functions with NQEs included.
As mentioned above, accurate PESs that incorporate long-range charge transfer interactions in condensed phases and interfaces are crucial for predictive spectral simulations. Traditional empirical force fields are computationally efficient, but they lack the accuracy and transferability required to capture the complex processes involved in condensed phases and interfaces. On the other hand, ab initio PIMD simulations where forces felt by each nucleus are calculated on-the-fly using electronic structure methods are prohibitively expensive for obtaining converged real-time dynamical properties, especially at low temperatures.16 Machine learning interatomic potentials (MLIPs) have recently been shown to be able to bridge the gap between the accuracy and the cost of these simulations.17 This is due to the ability of MLIPs in leveraging data-driven approaches to approximate PESs with near ab initio accuracy at a fraction of the computational expense.18 One of the early established MLIP architectures that uses artificial neural networks (ANNs) is the high-dimensional neural network potentials (HDNNPs), which are widely adopted to construct PESs for both gas-phase and condensed-phase molecular systems.19
@@
In this work, 4G-HDNNPs are developed and integrated with different path integral methods for the accurate prediction of the IR and vSGF spectra of representative bulk water and air-water interface systems. Water’s complex structure arises from its dynamic hydrogen-bond (H-bond) network, with differential intra- versus inter-molecular H-bonds, which significantly influence its vibrational spectral properties.26 Recent studies have shown that this complex network supports propagating optical phonon-like modes, indicating coherent long-range interactions.27 These long-range interactions play a crucial role in water’s vibrational dynamics, affecting energy transfer processes and the overall behavior of its H-bond network. To achieve fast and efficient spectral calculations, an active learning framework with query-by-committee (QbC) is developed and employed for the first time to facilitate training of the 4G-HDNNP models. Our QbC approach enables data-efficient training through an iterative selection of the most distinct and informative configurations. Thus, we refer to our potentials 4G-HDCNNPs, where ”C” stands for committee.
Although attempts have been made to combine HDNNPs with PIMD simulations before,28 this work is the first report to directly employ fourth-generation PI simulations for the calculations of IR and vSFG spectra, bypassing the need for the explicit training of the dipoles. The 4G-HDCNNPs reported here have two pivotal advances compared to the second-generation HDNNP (2G-HDNNP) models. First, 4G-HDCNNPs incorporate explicit non-local charge-transfer interactions, capturing interactions that 2G models neglect and thereby delivering higher accuracy in energies and forces for environments with significant charge transfer and polarization effects.22, 24, 29 Second, as the 4G architecture predicts atomic charges, charge-dependent observables such as dipole moment can be evaluated directly, without the approximations that 2G models require.
The rest of the paper is organized as follows: in Sec. 2 we provide an overview of the theoretical background of 4G-HDCNNPs. The details of model training and dynamical simulations in calculating IR and vSFG spectra are given in Sec. 3. The results and accompanying discussions are presented in Sec. 4. We end this work with concluding remarks and a discussion of future directions in Sec. 5.
@@ 2 Theory Details
4G-HDNNPs seamlessly combine atom-centered descriptors with environment-dependent charges obtained from a global charge equilibration, taking into account both local short-range interactions and non-local electrostatics. This is achieved by employing one set of neural networks that predict environment-dependent electronegativities for long-range electrostatics and a second set of neural networks that model the local short-range contributions. The schematic of the networks and their interrelation is depicted in Fig. 1. This enables the potential to respond to non-local charge transfer and polarization without sacrificing accuracy in capturing local interactions. The total energy in a 4G-HDNNP is, therefore, expressed as22
@@ 2.2 IR Spectral Simulations
All IR spectra are calculated from the Fourier transform of the total cell dipole-derivative autocorrelation function, following our previous works,15, 31
@@
2.3 vSFG Spectra for Probing Interfaces
vSFG is a second-order non-linear vibrational spectroscopic technique whose response is governed by the second‑order susceptibility, $\chi^{(2)}$.34 This highly sensitive technique is shown to be exceptionally useful in probing molecules at interfaces, as it yields no signal in centrosymmetric bulk media. The resonant component of the second-order vSFG response function can be approximately calculated using the surface-specific velocity-velocity correlation function (ssVVCF)35, 36
$\chi^{\text{res},(2)}_{aac}(\omega)=$ $\frac{Q(\omega)}{i\omega^{2}}\int_{0}^{\infty}dt\,e^{-i\omega t}$
$\times\left\langle\sum_{i,j}g_{t}(r_{ij}(0);r_{t})\,\dot{r}^{\text{OH}}_{c,i}(0)\,\frac{\dot{{r}}^{\text{OH}}_{j}(t)\cdot{r}^{\text{OH}}_{j}(t)}{|{r}^{\text{OH}}_{j}(t)|}\right\rangle.$ (10)
Here, $r_{ij}(t)$ is the distance between the center of masses of O-H groups $i$ and $j$ at time $t$, and $g_{t}(r_{ij}(0);r_{t})$ is a cutoff function that controls the cross-correlation within a specified radius $r_{t}$,
$g(r_{ij}(t);r_{t})=\begin{cases}1&r_{ij}(t)\leq r_{t}\\ 0&r_{ij}>r_{t}.\end{cases}$
The ${r}^{\text{OH}}_{j}(t)$ and $\dot{{r}}^{\text{OH}}_{j}(t)$ terms denote distances and velocities of the O-H bond vector $j$ at time $t$, respectively. The quantum correction factor $Q(\omega)$ is used for the classical spectra as 35
$Q(\omega)=\left(\frac{\hbar\omega/kT}{1-e^{-\hbar\omega/kT}}\right).$ (12)
The imaginary part of $\chi^{(2)}$ contains direct information on the vibrational resonances37, which is reported in this work.
@@ 3 Simulation Details
The bulk water simulation box contains 64 water molecules in a cubic box of size 12.42 Å, corresponding to a density of 0.997 g/cm3 at 298 K. For the interfacial simulations, 96 water molecules were placed in a 14.21 Å $\times$ 14.21 Å $\times$ 39.21 Å  rectangular prism with a 12.5 Å  vacuum layer on each side of the slab along the $z$-direction. Periodic boundary conditions were used in all three directions. The reference ab initio MD (AIMD) trajectories were generated employing the revPBE0-D3 dispersion corrected hybrid functional38, 39, 40, 41 in CP2K 42 using the Quickstep module43. Reference atomic charges were calculated using the DDEC6 charge partitions.44 Previous studies have shown that NQE-included hybrid revPBE0-D3 simulations produce accurate structures and vibrational spectra for bulk water.45, 16 The initial configuration for bulk water was generated using Packmol46 and equilibrated for 10 ps using classical revPBE-D3 AIMD simulations in the NVT ensemble with a CSVR thermostat and a 1 fs time step. This was followed by 50 ps revPBE0-D3 AIMD simulations, with the snapshots from the final 40 ps used for training.
All 4G-HDCNNP model trainings were performed using N2P2 (see Fig. 1).47 The radial and angular symmetry functions for O and H atoms were taken from Ref. 20 and are provided in SI Tables S1 and S2. Two hidden layers, each with 15 nodes, were used for both electrostatics and short-range neural networks. The active learning with committee neural network architecture, as implemented in an in-house modified version of AML 20, was employed for efficient sampling and accurate model generations. All active learning parameters were extensively benchmarked. It was found that a committee of four models ($n=4$) is sufficient for all charge, energy, and force trainings (see SI Fig. S1). Also, 15 epochs were found to be adequate for both the electrostatics and short-range neural networks. All 4 committee models use the same set of symmetry functions to represent the local environments, but differ in their initial random seed. These variations ensure a sufficiently diverse committee of NNPs. In our active learning workflow, we first selected 20 maximally distant configurations in the first iteration. For all subsequent iterations, 20 configurations with the highest force disagreement among the committee members were selected according to Eq. 1320, where $n$ is the committee index and $\alpha$ is the atom index.
@@
The Generation 1 models were trained on a set of bulk classical configurations with $n=4$ committee models in the QbC cycle. For this purpose, a candidate pool of 4000 distant and random structures was harvested from the revPBE0-D3 AIMD simulation. The selection loop ensured that no configuration was sampled more than once, and evaluated the entire candidate pool at each round. The subsequent generations were trained with both bulk and air-water interface configurations. Once Generation 1 models were trained, MD and PIMD simulations were performed using an in-house modified version of DL_POLY Quantum.48, 31. DL_POLY Quantum is a highly modular, sustainable, and scalable MD simulation platform for large-scale long-time classical and PI simulations of condensed phase and interfacial systems with the essential NQEs included. Generation 2 models were trained with the configuration pool generated from DL_POLY Quantum simulations for classical and partially quantum nuclei. At this stage, NQEs were gradually introduced with configurations from partially converged PIMD simulations with 8 beads. Both bulk and interfacial configurations were added to the Generation 2 pool. Similarly, Generation 3 was trained with configurations selected from fully converged PIMD simulations with 32 beads generated from both bulk and air-water interface simulations in DL_POLY Quantum. Similar to Ref. 20, a biasing potential was applied during the 4G-HDCNNP MD and PIMD simulations at every stage to stabilize the simulations and prevent the simulations from sampling regions of configuration space with large committee disagreements. Energy, force, and atomic charges for the selected configurations from Generations 2 and 3 were calculated using the same revPBE0-D3 hybrid dispersion-corrected functional and DDEC6 charge partition. All results reported here are obtained using the final Generation 3 models, unless stated otherwise.
3.2 Simulating IR and vSFG Spectra
@@
Using the trained final Generation 3 models, an initial 50 ps canonical (NVT) sampling was performed for bulk water using the PI Langevin equation (PILE) thermostat49 with timesteps of 1 fs and 0.5 fs for classical and PIMD simulations, respectively. After discarding equilibration configurations, 10 independent snapshots were extracted for microcanonical (NVE) simulations of 10 ps each with a timestep of 1 fs, 0.5 fs, and 0.25 fs for classical, TRPMD, and PA-CMD simulations, respectively. For each of the NVE trajectories, the initial 2 ps were discarded as equilibration, and the remaining 8 ps trajectory was used for spectral analysis. All dipole moments were calculated using the standard expression for a system of classical point charges: $\boldsymbol{\mu}=\sum_{i}Q_{i}\mathbf{R}_{i}$, where $Q_{i}$ is the atomic charge associated with atom $i$ and $\mathbf{R}_{i}$ is its position. Similar to Ref. 50, every water molecule was first reconstructed with minimum-image connectivity. The center of mass for each water molecule was then shifted to the origin, and the $x$, $y$, and $z$ components of $\boldsymbol{\mu}$} were evaluated. Using these dipoles directly, IR spectra were calculated from the Fourier transform of the total cell dipole-derivative autocorrelation functions with a Hann window cutoff of $\tau_{0}$ = 0.3 ps. The lineshape sensitivity to different time constants is demonstrated for TRPMD simulations in SI Fig. S8. The same converged NVE trajectories were used for vSFG spectral simulations. Similar to Ref. 51, a 4 Å-thick interfacial region on both sides of the water slab was selected to represent surface water molecules contributing to the vSFG signal. A 6 Å cutoff radius was used for all vSFG calculations to ensure the inclusion of intermolecular coupling effects.
@@ 4.1 Active Learning and Model Performance
To benchmark and evaluate our active learning workflow in Generation 1, a separate validation set consisting of 500 distant AIMD-generated bulk configurations was used. For n=2, 4, and 8 models, the charge and force errors on that validation set are illustrated in SI Fig. S1. The MAE values converge quickly with increasing the number of committees and plateau within 300 configurations. As such, Generation 1 was trained with 300 QbC selected configurations. Such choices and error levels are consistent with prior reports on water using active learning and 2G-HDNNP committee models, where a few hundred reference configurations were found to be sufficient for achieving high accuracy.20 The charge RMSE for this generation was calculated to be 5.97 me and 5.91 me for the training and test sets, respectively. The energy RMSE is found to be 0.45 meV/atom and 0.41 meV/atom, and the force RMSE was calculated to be 0.05 eV/Å for both the training and test sets. These values are consistent with those reported previously (the comparisons are provided in the SI Table S3).22, 24 Calculated correlations between the reference and 4G-HDCNNP predicted charges, energies, and forces for the training and test sets are given in SI Fig. S2.
Fig. 2 illustrates the calculated MAEs across all three generations.
@@
This evaluation was performed on another separate validation set of 2000 configurations from four different trajectories simulated with the final Generation 3 models. The first set of configurations included 500 distinct classical bulk simulation snapshots. The following 500 configurations were generated from a converged 32-bead PIMD simulation of bulk water. The 3rd and 4th sets of 500 configurations each originated from similar classical and PIMD simulations of interfacial water. All reference energy, force, and DDEC6 charges were calculated at the revPBE0-D3 level. As can be seen from Fig. 2, the classical force MAEs are calculated to be 49 meV/Å, 44 meV/Å, and 44 meV/Å  for Generation 1-3 models, respectively. For configurations with quantum nuclei, the MAEs are 86 meV/Å, 55 meV/Å, and 53 meV/Å  for Generation 1-3 models, respectively. The first-generation models trained with only classical configurations perform poorly in describing both bulk and interfacial configurations with quantum nuclei, which improves with subsequent generations as NQEs are introduced during training. All MD and PIMD production simulations using the final trained 4G-HDCNNPs were performed with $n=4$ committee models for Generations 1 and 2. To reduce the computational cost of our simulations without compromising accuracy, for Generation 3, we reduced the number of models to $n=2$, as it was found to yield similar results while being comparatively faster, especially for path integral simulations involving air-water interface systems. This enabled us to perform longer trajectories with no loss of accuracy. The comparison among n = 1, 2, 3, and 4 models, validated using bulk water radial distribution functions (RDFs) calculated using Generation 3 models, is shown in SI Fig. S6.

### v2

# Predictive quantum vibrational spectra through active learning 4G-NNPs

@@ Abstract
Predictive simulation of vibrational spectra of complex condensed-phase and interface systems with thousands of degrees of freedom has long been a challenging task of modern condensed matter theory. In this work, fourth-generation high-dimensional committee neural network potentials (4G-HDCNNPs) are developed using active learning and query-by-committee, and introduced to the essential nuclear quantum effects (NQEs) as well as conformational entropy and anharmonicities from path integral (PI) molecular dynamics simulations. Using representative bulk water and air-water interface test cases, we demonstrate the accuracy of the developed framework in infrared spectral simulations. Specifically, by seamlessly integrating non-local charge transfer effects from 4G-HDCNNPs with the NQEs from PI methods, our introduced methodology yields accurate infrared spectra using predicted charges from the 4G-HDCNNP architecture without explicit training of dipole moments. The framework introduced in this work is simple and general, offering a practical paradigm for predictive spectral simulations of complex condensed phases and interfaces, free from empirical parameterizations and ad hoc fitting.
@@ 1 Introduction
Vibrational spectroscopy is an excellent tool for probing complex processes across different condensed-phase and interface environments.1, 2, 3, 4 To obtain accurate predictive vibrational spectra, free from ad hoc fittings or empirical parameterizations, it has been well-established that robust dynamical methodologies that incorporate anharmonicities and conformational entropies, as well as the essential nuclear quantum effects (NQEs) are needed to be developed and integrated with accurate potential energy surfaces (PESs) that include non-local charge transfer interactions.5, 6 Path integral molecular dynamics (PIMD) offers an accurate and efficient route to include NQEs by representing a quantum particle as a series of identical classical particles, referred to as beads connected via harmonic springs to form a closed loop, referred to as a ring polymer.7, 8, 9 Accurate equilibrium properties that incorporate NQEs can be calculated using PI Monte Carlo (PIMC) or PI molecular dynamics (PIMD) simulations.8 Similarly, approximate real-time PI methods, including thermostatted ring polymer MD (TRPMD),10 partially adiabatic centroid MD (PA-CMD)11, and a myriad of methods branched from them,12, 13, 14, 15, can then be used for the simulation of dynamical properties including vibrational spectra through the calculations of appropriate time correlation functions with NQEs included.
As mentioned above, accurate PESs that incorporate long-range charge transfer interactions in condensed phases and interfaces are crucial for predictive spectral simulations. Traditional empirical force fields are computationally efficient, but they lack the accuracy and transferability required to capture the complex processes involved in condensed phases and interfaces. On the other hand, ab initio PIMD simulations where forces felt by nuclei are calculated on-the-fly using electronic structure methods are prohibitively expensive for obtaining converged real-time dynamical properties, especially at low temperatures.16 Machine learning interatomic potentials (MLIPs) have recently been shown to be able to bridge the gap between the accuracy and the cost of these simulations.17 This is due to the ability of MLIPs in leveraging data-driven approaches to approximate PESs with near ab initio accuracy at a fraction of the computational expense.18 One of the early established MLIP architectures that uses artificial neural networks (ANNs) is the high-dimensional neural network potentials (HDNNPs), which are widely adopted to construct PESs for both gas-phase and condensed-phase molecular systems.19
@@
In this work, 4G-HDNNPs are developed and integrated with different path integral methods for the accurate prediction of the IR spectra of representative bulk water and air-water interface systems. Water’s complex structure arises from its dynamic hydrogen-bond (H-bond) network, with differential intra- versus inter-molecular H-bonds, which significantly influence its vibrational spectral properties.26 Recent studies have shown that this complex network supports propagating optical phonon-like modes, indicating coherent long-range interactions.27 These long-range effects play a crucial role in water’s vibrational dynamics, affecting energy transfer processes and the overall behavior of its H-bond network. To achieve fast and efficient spectral calculations, an active learning framework using query-by-committee (QbC) is developed and employed to facilitate training of the 4G-HDNNP models. Our QbC approach enables data-efficient training through an iterative selection of the most distinct and informative configurations. Thus, we refer to our developed potentials 4G-HDCNNPs, where ”C” stands for committee.
Although attempts have been made to combine HDNNPs with PIMD simulations before,28 this work is the first report to directly employ fourth-generation PI simulations for the calculations of IR spectra, bypassing the need for the explicit training of the dipoles. The 4G-HDCNNPs reported here have two pivotal advances compared to the second-generation HDNNP (2G-HDNNP) models. First, 4G-HDCNNPs incorporate explicit non-local charge-transfer interactions, capturing interactions that 2G models neglect and thereby delivering higher accuracy in energies and forces for environments with significant charge transfer and polarization effects.22, 24, 29 Second, as the 4G architecture predicts atomic charges, charge-dependent observables such as dipole moment can be evaluated directly, without the approximations that 2G models require. Below, we first provide a brief theoretical background for 4G-HDCNNPs, along with details of model training and IR spectral simulations, followed by our simulated IR spectra for the considered bulk water and air-water interface systems.
@@ 2 Theory Details
4G-HDNNPs seamlessly combine atom-centered descriptors with environment-dependent charges obtained from a global charge equilibration, taking into account both local short-range interactions and non-local long-range electrostatics. This is achieved by employing one set of neural networks to predict environment-dependent electronegativities for long-range electrostatics and a second set of neural networks to model the local, short-range contributions. The schematic of the networks and their interrelation is depicted in Fig. 1. This enables the potential to respond to non-local charge transfer and polarization without sacrificing accuracy in capturing local interactions. The total energy in a 4G-HDNNP is, therefore, expressed as22
@@ 2.2 IR Spectral Simulations
Using 4G predicted charges directly and without explicit training of dipoles, IR spectra are calculated from the Fourier transform of the total cell dipole-derivative autocorrelation function, following our previous works,15, 31
@@ 3 Simulation Details
The bulk water simulation box contains 64 water molecules in a cubic box of size 12.42 Å, corresponding to a density of 0.997 g/cm3 at 298 K. For the interfacial simulations, 96 water molecules were placed in a 14.21 Å $\times$ 14.21 Å $\times$ 39.21 Å  rectangular prism with a 12.5 Å  vacuum layer on each side of the slab along the $z$-direction. Periodic boundary conditions were used in all three directions. The reference ab initio MD (AIMD) trajectories were generated employing the revPBE0-D3 dispersion corrected hybrid functional34, 35, 36, 37 in CP2K 38 using the Quickstep module39. Reference atomic charges were calculated using the DDEC6 charge partitions.40 Previous studies have shown that NQE-included hybrid revPBE0-D3 simulations produce accurate structures and vibrational spectra for bulk water.41, 16, 42 The initial configuration for bulk water was generated using Packmol43 and equilibrated for 10 ps using classical revPBE-D3 AIMD simulations in the NVT ensemble with a CSVR thermostat and a 1 fs time step. This was followed by 50 ps revPBE0-D3 AIMD simulations, with the snapshots from the final 40 ps used for training.
All 4G-HDCNNP model trainings were performed using n2p2 (see Fig. 1).44 The radial and angular symmetry functions for O and H atoms were taken from Ref. 20 and are provided in the Supporting Information Tables S1 and S2. Two hidden layers, each with 15 nodes, were used for both electrostatics and short-range neural networks. The active learning with committee neural network architecture, as implemented in an in-house modified version of AML 20, was employed for efficient sampling and accurate model generations. All active learning parameters were extensively benchmarked. It was found that a committee of four models ($n=4$) is sufficient for all charge, energy, and force trainings (see the Supporting Information Fig. S1). Also, 15 epochs were found to be adequate for both the electrostatics and short-range neural networks. All 4 committee models use the same set of symmetry functions to represent the local environments, but differ in their initial random seed. These variations ensure a sufficiently diverse committee of NNPs. In our active learning workflow, we first selected 20 maximally distant configurations in the first iteration. For all subsequent iterations, 20 configurations with the highest force disagreement among the committee members were selected according to Eq. 1020, where $n$ is the committee index and $\alpha$ is the atom index.
@@
The Generation 1 models were trained on a set of bulk classical configurations with $n=4$ committee models in the QbC cycle. For this purpose, a candidate pool of 4000 distant and random structures was harvested from the revPBE0-D3 AIMD simulation. The selection loop ensured that no configuration was sampled more than once, and evaluated the entire candidate pool at each round. The subsequent generations were trained with both bulk and air-water interface configurations. Once Generation 1 models were trained, MD and PIMD simulations were performed using an in-house modified version of our DL_POLY Quantum code.45, 31. DL_POLY Quantum is a highly modular, sustainable, and scalable MD simulation platform for large-scale long-time classical and PI simulations of condensed phase and interfacial systems with the essential NQEs included. Generation 2 models were trained with the configuration pool generated from DL_POLY Quantum simulations for classical and partially quantum nuclei. At this stage, NQEs were gradually introduced with configurations from partially converged PIMD simulations with 8 beads. Both bulk and interfacial configurations were added to the Generation 2 pool. Similarly, Generation 3 was trained with configurations selected from fully converged PIMD simulations with 32 beads generated from both bulk and air-water interface simulations in DL_POLY Quantum. Similar to Ref. 20, a biasing potential was applied during the 4G-HDCNNP MD and PIMD simulations at every stage to stabilize the simulations and prevent the simulations from sampling regions of configuration space with large committee disagreements. Energy, force, and atomic charges for the selected configurations from Generations 2 and 3 were calculated using the same revPBE0-D3 hybrid dispersion-corrected functional and DDEC6 charge partition. All results reported here are obtained using the final Generation 3 models, unless stated otherwise.
3.2 Simulating IR Spectra of Bulk and Interfacial Water
@@
Using the trained final Generation 3 models, an initial 50 ps canonical (NVT) sampling was performed for bulk water using the PI Langevin equation (PILE) thermostat46 with timesteps of 1 fs and 0.5 fs for classical and PIMD simulations, respectively. After discarding equilibration configurations, 10 independent snapshots were extracted for microcanonical (NVE) simulations of 10 ps each with a timestep of 1 fs, 0.5 fs, and 0.25 fs for classical, TRPMD, and PA-CMD simulations, respectively. For each of the NVE trajectories, the initial 2 ps were discarded as equilibration, and the remaining 8 ps trajectory was used for spectral analysis. All dipole moments were calculated using the standard expression for a system of classical point charges: $\boldsymbol{\mu}=\sum_{i}Q_{i}\mathbf{R}_{i}$, where $Q_{i}$ is the atomic charge associated with atom $i$ and $\mathbf{R}_{i}$ is its position. Similar to Ref. 47, every water molecule was first reconstructed with minimum-image connectivity. The center of mass for each water molecule was then shifted to the origin, and the $x$, $y$, and $z$ components of $\boldsymbol{\mu}$} were evaluated. Using these dipoles directly, IR spectra were calculated from the Fourier transform of the total cell dipole-derivative autocorrelation functions with a Hann window cutoff of $\tau_{0}$ = 0.3 ps. The lineshape sensitivity to different time constants is demonstrated for TRPMD simulations in the Supporting Information Fig. S8.
@@ 4.1 Active Learning and Model Performance
To benchmark and evaluate our active learning workflow in Generation 1, a separate validation set consisting of 500 distant AIMD-generated bulk configurations was used. For n=2, 4, and 8 models, the charge and force errors on that validation set are illustrated in the Supporting Information Fig. S1. The MAE values converge quickly with increasing the number of committees and plateau within 300 configurations. As such, Generation 1 was trained with 300 QbC selected configurations. Such choices and error levels are consistent with prior reports on water using active learning and 2G-HDNNP committee models, where a few hundred reference configurations were found to be sufficient for achieving high accuracy.20 The charge RMSE for this generation was calculated to be 5.97 me and 5.91 me for the training and test sets, respectively. The energy RMSE is found to be 0.45 meV/atom and 0.41 meV/atom, and the force RMSE was calculated to be 0.05 eV/Å for both the training and test sets. These values are consistent with those reported previously (the comparisons are provided in the Supporting Information Table S3).22, 24 Calculated correlations between the reference and 4G-HDCNNP predicted charges, energies, and forces for the training and test sets are given in the Supporting Information Fig. S2.
@@
Fig. 2 illustrates the calculated MAEs across all three generations. This evaluation was performed on another separate validation set of 2000 configurations from four different trajectories simulated with the final Generation 3 models. The first set of configurations included 500 distinct classical bulk simulation snapshots. The following 500 configurations were generated from a converged 32-bead PIMD simulation of bulk water. The 3rd and 4th sets of 500 configurations each originated from similar classical and PIMD simulations of interfacial water. All reference energy, force, and DDEC6 charges were calculated at the revPBE0-D3 level. As can be seen from Fig. 2, the classical force MAEs are calculated to be 49 meV/Å, 44 meV/Å, and 44 meV/Å  for Generation 1-3 models, respectively. For configurations with quantum nuclei, the MAEs are 86 meV/Å, 55 meV/Å, and 53 meV/Å  for Generation 1-3 models, respectively. Expectedly, the first-generation models trained with only classical configurations perform poorly in describing both bulk and interfacial configurations with quantum nuclei, which improves with subsequent generations as NQEs are introduced during training. All MD and PIMD production simulations using the final trained 4G-HDCNNPs were performed with $n=4$ committee models for Generations 1 and 2. To reduce the computational cost of our simulations without compromising accuracy, for Generation 3, we reduced the number of models to $n=2$, as it was found to yield similar results while being comparatively faster, especially for path integral simulations involving air-water interface systems. This enabled us to perform longer trajectories with no loss of accuracy. The comparison among n = 1, 2, 3, and 4 models, validated using bulk water radial distribution functions (RDFs) calculated with Generation 3 models, is presented in the Supporting Information Fig. S6.